OpenStudy (anonymous):

The rate at which a plant grows is constant. At the 3rd week it is 12 inches tall and at the 5th week it is 20 inches tall. If this information is plotted on a coordinate plane, what is the y-coordinate of the the point (1, y)? A) 1 B) 2 C) 4 D) 9

OpenStudy (anonymous):

@paki or anybody else help me

OpenStudy (unklerhaukus):

find the rate of growth by taking the relative difference between the know points

OpenStudy (unklerhaukus):

\[\frac{\mathrm dy}{\mathrm dx} = m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

if the rate grow is constant , this function is takes the number of week and give you the tall of the planet. f(x) =y , x number of weeks and y is tall of planet. given :- y(3) =12cm and y(5) = 20cm

OpenStudy (anonymous):

to know the constant rate of growing we must to know how many growing in one week or to know the slope of the function :- slope m = (y1-y2)/(x1-x2)

OpenStudy (anonymous):

(5,20) and (3,12); m =(20-12)/(5-3) = 4; this mean the planet is growing 4cm every week; functions is y =4x; y(1) = 4cm this answer of the question